Find the magnitude of the following vector v= 5i - 2j
Found 2 solutions by stanbon, opeodu:
Answer by stanbon(75887)

Find the magnitude of the following vector v= 5i - 2j
---
r = sqrt[5^2+(-2)^2]
---
r = sqrt[25+4]
r = sqrt(29)
=================
Cheers,
Stan H.

Answer by opeodu(24)

|v|=(5^2+(-2)^2)^1/2 =(25+4)^1/2 =29^1/2